OpenStudy (anonymous):

Rationalize the denominator.. \[\frac{ -11x^3 }{ \sqrt{7x+1} }\] @Loser66

OpenStudy (opcode):

Well if you only want a rational denominator bring the \(\sqrt{}\) to the top. \[\dfrac{ -11x^3 \times \sqrt{7x+1}}{\sqrt{7x+1} \times \sqrt{7x+1}}\] Which becomes: \[\dfrac{ -11x^3 \times \sqrt{7x+1}}{7x+1}\] Khan academy has a topic about rationalizing the denominator: https://www.khanacademy.org/math/algebra/exponent-equations/simplifying-radical-expressions/v/how-to-rationalize-a-denominator

OpenStudy (anonymous):

Alright, thanks bro. Sorry for late response, was eating :). Appreciated very much mate.

OpenStudy (opcode):

No problems, mate. Glad I was of any help :-).